OpenStudy (anonymous):

Points V, W, X, Y, and Z are graphed on a number line. The coordinate of point V is –8. What points are at a distance of 5 units from point V? Choose all answers that are correct. A. W = –13 B. X = –8 C. Y = –3 D. Z = 3

OpenStudy (anonymous):

@k12andstudyislandhelp

OpenStudy (anonymous):

Points Y and Z are the same distance from point X on a number line. The coordinate of point X is –7 and the coordinate of point Y is –1. What is the coordinate of point Z?

OpenStudy (anonymous):

both of them

OpenStudy (anonymous):

ok, i dont think i know this. Sorry. I really am.

OpenStudy (anonymous):

that is ok thanks

OpenStudy (anonymous):

@undeadknight26 @King.Void.

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (anonymous):

@Somy

OpenStudy (sleepyhead314):

-8 - 5 = -13 -8 + 5 = -3